Next Question Try another similar question , reattempt last question below, or select another question. Given the numbers in the set { ,1, 3,25가 list the irrational mumbers. -9,-5,-0.111 7 n 8 on 9 ion 10 tion 11 stion 12 estlon 13 25 TIP uestion 14 Select all correct answers Get Help Video eBook Points possible: 1 Question 15 de: 1.4/15 nt Version nlimited attempts の
Answer #1

Irrational numbers are that which we can not express as a fraction.

-9 we can write as -9 = \frac{-9}{1}

-5 can write as: -5 = \frac{-5}{1}

-0.1111........... can write as -0.1111...........= \frac{-1}{9}

1 can write as 1 = \frac{1}{1}

3 can write as 3 = \frac{3}{1}

\frac{25}{2} is already in fractions.

So these all are rational numbers.

Only \pi is irrational number.